About 4.9 million households had one brand of personal computers in 2001. The use of these computers grew at an average rate of .275 million households a year. In 2001, about 2.5 millions households used another type of computer. The use of these computers grew at an average rate of .7 million households a year. How long will it take for the two types of computers to be in the same number of households?
Answer by stanbon(75887)

About 4.9 million households had one brand of personal computers in 2001. The use of these computers grew at an average rate of .275 million households a year.
Eq: N = 4.9 + 0.275x million
----------------
In 2001, about 2.5 millions households used another type of computer. The use of these computers grew at an average rate of .7 million households a year.
Eq: P = 2.5 + 0.7x million
---------------------------------
How long will it take for the two types of computers to be in the same number of households?
4.9 + 0.275 = 2.5 + 0.7x
2.4 = 0.425x
x = 5.647 years
=================
Cheers,
Stan H.